# Stripe dark solitons in a nonlinear optical fiber with second- and fourth-order dispersion

### 11 Dec 2023 v1

@@ Abstract
We study the soliton excitations in a nonlinear optical fiber with the second- and fourth-order dispersion, and find the emergence of stripe dark solitons (SDSs) in addition to the existing stripe bright solitons (SBSs). The SDSs can exhibit the fundamental structures and their counterparts without a phase step, as well as the compound structures with different numbers of humps. By the modified linear stability analysis, we regard the formation of fundamental SDS as the competition between its periodicity and localization, and analytically give its existence condition, oscillation frequency, and transmission stability, which show great agreements with numerical results. We also show good predictions for the characteristics of SBS, and find that SBS will become the pure-quartic soliton when its periodicity and localization keep balance. Our results can be observable with current experimental techniques in nonlinear optical fibers, and our method can also guide the discovery of stripe solitons in other physical systems.
@@ I Introduction
Periodicity and localization are two fundamental but distinct characteristics of nonlinear waves. The periodicity is a commonality of all waves and dominates various wave phenomena Hecht-book ; Cronin-2009 , while the (self-induced) localization dominates the formation of nonlinear localized waves Mollenauer-1980 ; Weiner-1988 ; Onorato-2013 ; Dudley-2014 . As a kind of localized wave broadly existing in nonlinear physical systems, soliton has been attracting lots of attentions and has wide applications by virtue of its property of stable propagation Kivshar-book ; Agrawal-book ; Kevrekidis-book ; Haus-1996 ; McDonald-2014 . Soliton can produce periodicity-dominated phenomena like interference McDonald-2014 ; Helm-2015 ; Zhao-2016 and diffraction Xiong-2023 ; Gao-2023 , and it also has localization-dominated phenomena like self-bound propagation and elastic collision Zabusky-1965 ; Martin-2007 . In the previous studies, the periodicity and localization of a soliton were always treated separately, and little attention has been paid to their combined effects. However, in nonlinear fibers with the fourth-order dispersion (FOD), the bright solitons with oscillating tails Akhmediev-1994 ; Buryak-1995 ; Blanco-2016 ; Tam-2019 and even stripe structures Zakharov-1998 ; Tam-2020 were discovered, which are the manifestation of the competition between periodicity and localization. Therefore, to figure out the formation mechanism of these solitons and find more kinds of solitons in this system, it is necessary to quantify the two characteristics and analyze the competition between them.
In this paper, we study the soliton excitations in a nonlinear optical fiber with the second-order dispersion and FOD. We find that stripe dark solitons (SDSs) exist in this system and have more interesting structures than the stripe bright solitons (SBSs) shown in Refs. Zakharov-1998 ; Tam-2020 . These structures contains the fundamental SDSs and their counterparts without a phase step, as well as the compound structures with different numbers of humps. By the modified linear stability analysis (MLSA) method Gao-2020 ; Gao-2021 , a fundamental SDS is considered as the result of the competition between periodicity and localization, and furthermore we quantitatively give its existence condition and oscillation frequency. Based on the derived oscillation frequency, we analyze the modulation instability (MI) of SDSs and give its stability condition. Besides, we apply the MLSA to the SBSs to give a more accurate prediction for their oscillation frequency, and find that the pure-quartic soliton is a special kind of SBS when its periodicity and localization keep balance.
II Model and the existence condition of stripe solitons
@@
where $\beta_{2}=\beta^{\prime}_{2}U_{z}/U_{t}^{2}$, $\beta_{4}=\beta^{\prime}_{4}U_{z}/U_{t}^{4}$, and $P_{0}$ is the background value of the incident light power. The two quantities, $U_{z}$ and $U_{t}$, respectively have the units of space and time. In Ref. Runge-2020 , to generate a pure-quartic soliton in a fiber laser, the parameters were set as $\beta^{\prime}_{2}=21.4\,{\rm ps^{2}/km}$, $\beta^{\prime}_{4}=-80\,{\rm ps^{4}/km}$, $\gamma=1.6\,{\rm/W/km}$, and $P_{0}=0.37\,{\rm W}$. When $U_{z}=1/\gamma P_{0}=1.69\,{\rm km}$ and $U_{t}=6\,{\rm ps}$, one can obtain $\beta_{2}=1$ and $\beta_{4}=-0.13$. In our work, the range of $\beta_{2}$ and $\beta_{4}$ will be extended to generate more kinds of solitons.
Figure 1: (Color online) Existence condition of solitons under different second-order dispersion $\beta_{2}$ and FOD $\beta_{4}$. They contain the regions of SDS (stripe dark soliton), DS (the traditional dark soliton), SBS (the stripe bright soliton), and BS (the traditional bright soliton). The boundary of SDS region has the expression $\beta_{4}=3\beta_{2}^{2}/4a_{0}^{2}$, and the boundaries of SBS region is $\beta_{4}=-3\beta_{2}^{2}/2\beta$, where $a_{0}$ and $\beta$ are respectively the SDS’s background amplitude and the SBS’s propagation constant. The color scale describes the strength $\varphi$ of a soliton’s periodicity relative to its localization in the range from 0 to $\pi/2$. The black dashed line denotes the case of $\beta_{2}=0$, where a pure-quartic soliton can exist. The parameters are set as $a_{0}=\beta=1$.
The existence condition of solitons in the model (2) is shown in Fig. 1. Different kinds of solitons correspond to different regions in the $\beta_{2}$-$\beta_{4}$ plane. The details are as follows.
@@
  • When $\beta_{4}=0$, the model (2) becomes the standard nonlinear Schödinger model, where only the traditional dark ($\beta_{2}>0$) or bright ($\beta_{2}<0$) solitons can exist.
@@
  • When $\beta_{4}>{3\beta_{2}^{2}}/{4a_{0}^{2}}$, the SDSs exist and their periodicity gets stronger with $\beta_{2}$ decreasing [see Eq. (15)].
@@
  • When $\beta_{4}\leq{3\beta_{2}^{2}}/{4a_{0}^{2}}$ and $\beta_{2}>0$, the traditional dark soliton exists, which has no stripe.
@@
  • When $\beta_{4}<-{3\beta_{2}^{2}}/{2\beta}$, the SBSs exist and their periodicity gets stronger with $\beta_{2}$ increasing [see Eq. (25)]. In particular, it becomes the so-called pure-quartic soliton when $\beta_{2}=0$ Blanco-2016 .
@@
  • When $\beta_{4}<-{3\beta_{2}^{2}}/{2\beta}$ and $\beta_{2}<0$, the traditional bright soliton exists, which has no stripe.
@@
In the above conditions, $a_{0}(>0)$ denotes the plane-wave background’s amplitude of dark solitons, and $\beta(>0)$ denotes the propagation constant of bright solitons. In other regions of the $\beta_{2}$-$\beta_{4}$ plane, no soliton exists. We define the quantity $\varphi$ by $\varphi=\tan^{-1}({\omega_{s}}/{\eta_{s}})$ in the range from 00 to $\pi/2$ to describe the strength of a wave’s stripe characteristic. $\omega_{s}$ and $\eta_{s}$ are respectively the parameters determining the periodicity and localization of a wave, whose expressions can be seen in Eqs. (III) and (IV) for SDS and SBS, respectively. With $\varphi$ increasing, the wave’s periodicity gets stronger and the wave’s localization gets weaker. The stripe solitons with different shapes can be regarded as the results of the competition between periodicity and localization in different proportions. In particular, for a pure-quartic soliton Blanco-2016 , the two parameters $\omega_{s}$ and $\eta_{s}$ are equal, which indicates that its periodicity and localization keep balance.
@@ III SDS
In the model (2), we find two major kinds of SDSs, which severally have the fundamental and compound structures. At first, let us focus on the SDSs with fundamental structures and their counterparts without a phase step. Considering that it is quite difficult to analytically solve Eq. (2), we numerically solve it by the biconjugate gradient method Press-book . This numerical method has been successfully applied to obtain dark solitons or vertices in different models of Bose-Einstein condensates Winiecki-1999 ; Edmonds-2016 . By this method, when $\beta_{4}=1$, we obtain the numerical solutions of dark solitons under different $\beta_{2}$ and show them in Fig. 2. When $\beta_{2}=-1$, we find two kinds of fundamental SDSs, which has or does not have a phase step (i.e., the key characteristic of dark solitons), and we show them at the top or bottom of Fig. 2 (a), respectively. Both of them have some stripes and oscillating tails, but their structures are different. Later, we set their wave functions as the initial states to numerically integrate the model (2), by the split-step Fourier method Yang-book . The evolution results are shown as the insets in Fig. 2 (a). The SDS with the phase step stably propagates for seven unit distance. At its final evolution stage, some new waves can be seen, which is generally induced by the modulation instability (MI). The similar result is also obtained for the SDS without a phase step, which propagates stably for six unit distance.
Figure 2: (Color online) Distributions of the wave functions $\psi$ of dark solitons (red dashed curves) and their amplitude $|\psi|$ (blue solid curves), when the second-order dispersion $\beta_{2}$ is (a) $-1$, (b) $-0.8$, (c) $-0.3$, (d) 0, and (e) 1.5. The upper and lower rows are the dark solitons with and without a phase step, respectively. The insets are their corresponding amplitude evolution plots. Other parameters are set as $\beta_{4}=1$ and $a_{0}=1$.
@@
For some larger values of $\beta_{2}$, the results are shown in Figs. 2 (b), (c), (d), and (e). The SDS has less stripes but can propagate for a longer distance, especially when $\beta_{2}>0$. Note that the SDS without a phase step can be regarded as two separated dark solitons when $\beta_{2}=1.5$, as shown in Fig. 2 (e). It indicates that, when the SDS dose not exist, there will not exist the dark soliton without a phase step.
@@
As shown in Fig. 2, the SDSs have both of periodicity and localization. For such a nonlinear wave, the MLSA method can quantitatively give its dynamical characteristics and guide its generation Gao-2020 ; Gao-2021 . Thus, we will apply this method to analyze the existence condition and characteristics of fundamental SDSs. It starts with the ansatz solution of a perturbed plane wave,
@@
where $\psi_{0}=a_{0}e^{ia_{0}^{2}z}$ is a plane wave with the amplitude $a_{0}$, and $u$ denotes a perturbation exerted on it. Substituting Eq. (3) into the model (2) and linearizing it about $u$, we can obtain
$i\frac{\partial u}{\partial z}-\frac{\beta_{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}+\frac{\beta_{4}}{24}\frac{\partial^{4}u}{\partial t^{4}}+a_{0}^{2}(u+u^{*})=0,$ (4)
where $u^{*}$ denotes the complex conjugation of $u$. Then, we assume the perturbation as the linear superposition of two conjugating components,
@@
where $A$ and $B$ are the complex amplitude of the two components, and $p(t,z)$ is a complex function. Substituting Eq. (5) into the perturbation’s model (4) and only considering the parts of $e^{p(t,z)}$, one can obtain two linear equations about $A$ and $B^{*}$. The condition that the equations have non-zero solutions is
$p_{z}=$ $\sqrt{M(2a_{0}^{2}-M)},$ (6)
@@
In the above expressions, $p_{t}$ and $p_{z}$ denote the partial derivative of $p(t,z)$ about $t$ and $z$, respectively.
@@
By this limit, the function $p_{t}(t,z)$ can be successfully transformed into a complex constant, and therefore offers great convenience for our analysis. Moreover, this limit can be conveniently applied to the function $p_{z}$ and produce
$p_{z}^{(-)}=$ $\sqrt{M^{(-)}(a_{0}^{2}-M^{(-)})},$ (10)
@@
It describes the propagating characteristics of the perturbation. To be specific, the perturbation’s propagation constant and growing rate relative to the plane wave are respectively
@@
Figure 3: (Color online) (a) Phase of the propagating function $p_{z}(t,z)$ under different $\omega$ and $\eta$, when $\beta_{2}=-0.5$, $\beta_{4}=1$, and $a_{0}=1$. The black solid and dashed curves denote $V_{\rm en}=0$ and $V_{\rm st}=0$, respectively. The white point with the coordinate $(\omega_{s},\eta_{s})$ is an exceptional point of $\arg(p_{z})$ function. The yellow curve with an arrow is the trajectory of this point with $\beta_{2}$ increasing. (b) Dependence of SDS’s oscillation frequency $\omega_{s}$ on $\beta_{2}$, when $\beta_{4}=1$. The red dots and the black curve are respectively the results from numerical evolution and the MLSA method. (c) Dependence of the gain value $G$ on $\omega$, when $\beta_{2}=-1$ (upper) or $\beta_{2}=-0.3$ (lower). The strong or weak MI is defined by whether $\omega_{s}$ locates in the MI band or not. (d) MI and MS distribution on the $\beta_{2}$-$\beta_{4}$ plane. The color scale describes the value of $g(\omega_{s})$.
From the insets of Fig. 2, we know that a SDS has a zero-velocity localized envelope and some zero-velocity stripes, which indicates $V_{\rm en}=0$ and $V_{\rm st}=0$. In Fig. 3 (a), when $\beta_{2}=-0.5$, $\beta_{4}=1$, and $a_{0}=1$, we show the curves of $V_{\rm en}=0$ and $V_{\rm st}=0$ in the $\omega$-$\eta$ plane. The coordinate of their intersection point $(\omega_{s},\eta_{s})$ is just the oscillation frequency and steepness of SDS, and the expressions are
$\omega_{s}=$ ${\rm Re}[\sqrt{\frac{-3\beta_{2}+2\sqrt{3}a_{0}\sqrt{\beta_{4}}}{\beta_{4}}}],$
$\eta_{s}=$ ${\rm Re}[\sqrt{\frac{3\beta_{2}+2\sqrt{3}a_{0}\sqrt{\beta_{4}}}{\beta_{4}}}].$ (14)
@@
$\omega_{s}^{2}+\eta_{s}^{2}=\frac{4\sqrt{3}a_{0}}{\sqrt{\beta_{4}}},\quad\omega_{s}^{2}-\eta_{s}^{2}=-\frac{6\beta_{2}}{\beta_{4}}.$ (16)
@@
The former is independent of $\beta_{2}$, and the latter is independent of $a_{0}$. Thus, the point $(\omega_{s},\eta_{s})$ always locates on a circle when $\beta_{4}$ and $a_{0}$ are fixed, and it moves towards the positive axis of $\eta$ with $\beta_{2}$ increasing, as shown in Fig. 3 (a). More interestingly, this point is just an exceptional point of the phase of $p_{z}$, and its winding number is $\pi$. However, for the relationship between SDS and exceptional points, a deeper understanding needs further investigations. Next, to verify the effectiveness of our prediction, the dependence of $\omega_{s}$ on $\beta_{2}$ is studied. When $\beta_{4}=1$ and $a_{0}=1$, we numerically measure the oscillation frequency of SDSs under different values of $\beta_{2}$ and show them by the red dots in Fig. 3 (b). One can see that $\omega_{s}$ gradually decreases into 0 with $\beta_{2}$ increasing. To obtain them, we measure the time interval between the first and second peaks on the right side of SDS as its time period $T_{s}$, and then the oscillation frequency can be calculated by $\omega_{s}=2\pi/T_{s}$. On the other hand, we also show the analytic prediction from the MLSA method by the black curve. There are great agreements between the two results.
@@
Figure 4: (Color online) The SDSs with compound structures. The upper row is the distribution of their wave functions $\psi$ (red dashed curves) and amplitude $|\psi|$ (blue solid curves), when $\beta_{2}=0$, $\beta_{4}=1$, and $a_{0}=1$. The lower row is their corresponding amplitude evolution.

### 07 Mar 2024 v2

# Dark solitons and their bound states in a nonlinear fiber with second- and fourth-order dispersion

@@ Abstract
We study the excitations of dark solitons in a nonlinear optical fiber with the second- and fourth-order dispersion, and find the emergence of striped dark solitons (SDSs) and some multi-dark-soliton bound states. The SDSs can exhibit time-domain oscillating structures on a plane wave, and they have two types: the ones with or without the total phase step, while the multi-dark-soliton bound states exhibit different numbers of amplitude humps. By the modified linear stability analysis, we regard the SDSs as the results of the competition between periodicity and localization, and analytically give their existence condition, oscillation frequency, and propagation stability, which show good agreements with numerical results. We also provide a possible interpretation of the formation of the existing striped bright solitons (SBSs), and find that SBS will become the pure-quartic soliton when its periodicity and localization keep balance. Our results provide the theoretical support for the experimental observation of striped solitons in nonlinear fibers, and our method can also guide the discovery of striped solitons in other physical systems.
@@ I Introduction
Periodicity and localization are two fundamental but distinct characteristics of nonlinear waves. The periodicity is a commonality of all waves and dominates various wave phenomena Hecht-book ; Cronin-2009 , while the (self-induced) localization dominates the formation of nonlinear localized waves Kivshar-book ; Onorato-2013 ; Dudley-2014 . As a widespread kind of nonlinear localized wave, soliton has been attracting lots of attentions and has wide applications by virtue of its property of stable propagation Mollenauer-1980 ; Weiner-1988 ; Agrawal-book ; Kevrekidis-book ; Haus-1996 ; McDonald-2014 . Soliton can produce periodicity-dominated phenomena like interference Snyder-1997 ; Kumar-2009 ; Zhao-2016 and diffraction Xiong-2023 ; Gao-2023 , and it also has localization-dominated phenomena like self-bound propagation and elastic collision Zabusky-1965 ; Stegeman-1999 . In the previous studies, the periodicity and localization of a soliton were always treated separately, and little attention has been paid to their combined effects. However, in nonlinear fibers with the fourth-order dispersion (FOD), the bright solitons with oscillating tails Akhmediev-1994 ; Buryak-1995 ; Blanco-2016 ; Tam-2019 and even stripe structures Zakharov-1998 ; Tam-2020 were discovered. They exhibit both of noticeable periodicity and localization, and therefore can be considered as the results of competition between the two characteristics. As we know, the modified linear stability analysis (MLSA) method has been used to predict the quantitative dynamics of nonlinear waves, including their periodicity and localization Gao-2020 ; Gao-2021 . Therefore, it provides the possibility to figure out the formation mechanism of striped solitons under the FOD and furthermore find more kinds of solitons in this system.
In this paper, we study the excitations of dark solitons in a nonlinear optical fiber with the second-order dispersion and FOD. We find that striped dark solitons (SDSs) and their bound states exist in this system, which exhibit more kinds of structures than the striped bright solitons Zakharov-1998 ; Tam-2020 and the multi-soliton bound states Buryak-1995 . The single-soliton state of SDS contains two types, i.e., the ones with or without the total phase step, while the multi-dark-soliton bound states have different numbers of amplitude humps in the time-domain distribution. By the MLSA method, a SDS is considered as the result of the competition between periodicity and localization, and furthermore we quantitatively give its existence condition and oscillation frequency. Based on the derived oscillation frequency, we analyze the modulation instability (MI) of plane-wave background of SDS and give its stability condition. Besides, we apply the MLSA to the SBSs to provide a possible interpretation of their formation, and find that the pure-quartic soliton is a special kind of SBS when its periodicity and localization keep balance.
II Model and the existence condition of striped solitons
@@
where $\beta_{2}=\beta^{\prime}_{2}U_{z}/U_{t}^{2}$, $\beta_{4}=\beta^{\prime}_{4}U_{z}/U_{t}^{4}$, and $P_{0}$ is the background value of the incident light power. The two quantities, $U_{z}$ and $U_{t}$, respectively have the units of space and time. In Ref. Runge-2020 , to generate a pure-quartic soliton in a fiber laser, the parameters were set as $\beta^{\prime}_{2}=21.4\,{\rm ps^{2}/km}$, $\beta^{\prime}_{4}=-80\,{\rm ps^{4}/km}$, $\gamma=1.6\,{\rm/W/km}$, and $P_{0}=0.37\,{\rm W}$. When setting $U_{z}=1/\gamma P_{0}=1.69\,{\rm km}$ and $U_{t}=6\,{\rm ps}$, one can obtain the dimensionless parameters, $\beta_{2}=1$ and $\beta_{4}=-0.13$. In our work, the range of $\beta_{2}$ and $\beta_{4}$ will be extended to generate more kinds of solitons.
In the model (2), some solitons and multi-soliton states have been studied Karlsson-1994 ; Piche-1996 ; Akhmediev-1994 ; Buryak-1995 ; Blanco-2016 ; Tam-2019 ; Zakharov-1998 ; Tam-2020 , including their existence conditions, stability, and even analytical solutions. All of these solitons belong to bright solitons, which have no plane waves as their background waves. In our work, the solitons we mainly concern about are the dark solitons, which exist on the plane-wave background, and therefore are essentially different from the bright solitons. According to our analysis, the existence condition of solitons in the model (2) is shown in Fig. 1. The solitons with different structures correspond to different regions in the $\beta_{2}$-$\beta_{4}$ plane. The details are as follows.
@@
  • The horizontal line: when $\beta_{4}=0$, the model (2) becomes the standard nonlinear Schödinger model, where only the traditional dark ($\beta_{2}>0$) or bright ($\beta_{2}<0$) solitons can exist.
@@
  • The region of SDS: when $\beta_{4}>{3\beta_{2}^{2}}/{4\beta}$, the SDSs exist and their periodicity gets stronger with $\beta_{2}$ decreasing [see Eq. (23)].
@@
  • The region of DS: when $\beta_{4}\leq{3\beta_{2}^{2}}/{4\beta}$ and $\beta_{2}>0$, the traditional dark soliton exists, which has no stripe.
@@
  • The region of SBS: when $\beta_{4}<-{3\beta_{2}^{2}}/{2\beta}$, the SBSs exist and their periodicity gets stronger with $\beta_{2}$ increasing [see Eq. (33)]. The solitons presented in Refs. Akhmediev-1994 ; Buryak-1995 ; Blanco-2016 ; Tam-2019 ; Zakharov-1998 ; Tam-2020 are in this region. In particular, it becomes the so-called pure-quartic soliton when $\beta_{2}=0$ Blanco-2016 .
@@
  • The region of BS: when $\beta_{4}<-{3\beta_{2}^{2}}/{2\beta}$ and $\beta_{2}<0$, the traditional bright soliton exists, which has no stripe. The solitons presented in Refs. Karlsson-1994 ; Piche-1996 are in this region.
@@
In the above conditions, $\beta(>0)$ denotes the propagation constant of solitons. In other regions of the $\beta_{2}$-$\beta_{4}$ plane, no soliton exists. We define the quantity $\varphi$ by $\varphi=\tan^{-1}({\omega_{s}}/{\eta_{s}})$ in the range from 00 to $\pi/2$ to describe the strength of a wave’s stripe characteristic. $\omega_{s}$ and $\eta_{s}$ are respectively the parameters determining the periodicity and localization of a wave, whose expressions can be seen in Eqs. (22) and (32) for SDS and SBS, respectively. With $\varphi$ increasing, the wave’s periodicity gets stronger and the wave’s localization gets weaker. The striped solitons with different shapes can be regarded as the results of the competition between periodicity and localization in different proportions. In particular, for a pure-quartic soliton Blanco-2016 , the two parameters $\omega_{s}$ and $\eta_{s}$ are equal, which indicates that its periodicity and localization keep balance.
@@
Figure 1: (Color online) Existence condition of solitons under different second-order dispersion $\beta_{2}$ and FOD $\beta_{4}$. They contain the regions of SDS (striped dark soliton), DS (the traditional dark soliton), SBS (the striped bright soliton), and BS (the traditional bright soliton). The boundary of SDS region has the expression $\beta_{4}=3\beta_{2}^{2}/4\beta$, and the boundary of SBS region is $\beta_{4}=-3\beta_{2}^{2}/2\beta$, where $\beta$ is the propagation constant of solitons. The color scale describes the strength $\varphi$ of a soliton’s periodicity relative to its localization in the range from 0 to $\pi/2$. The black dashed line denotes the case of $\beta_{2}=0$, where a pure-quartic soliton can exist. In all cases, the propagation constant is set as $\beta=1$.
Figure 2: (Color online) The dark solitons with (upper row) or without (lower row) the total phase step. Distributions of the wave functions $\psi$ of dark solitons (red dashed curves) and their amplitude $|\psi|$ (blue solid curves), when the second-order dispersion $\beta_{2}$ is (a) $-1$, (b) $-0.8$, (c) $-0.3$, (d) 0, and (e) 1.5. The insets are their corresponding amplitude evolution plots, where the horizontal lines are the spatial dividers between the solitons and the spontaneous oscillations induced by MI. Other parameters are set as $\beta_{4}=1$ and $\beta=1$.
@@ III SDS
III.1 Single-soliton state
At first, let us focus on the single-soliton states of SDSs and their counterparts without the total phase step. The term “the total phase step” is used to describe the phase difference of plane wave between the two sides of dark solitons, which is the key characteristic of dark solitons. It is known that the model (2) has the plane-wave solution,
$\psi_{0}=a_{0}e^{i\beta z},$ (3)
@@ III.1 Single-soliton state
where the amplitude $a_{0}$ and the propagation constant $\beta$ have the relationship $\beta=a_{0}^{2}$. In this section, we set $a_{0}=1$ so that $\beta=1$. Considering that it is quite difficult to analytically solve Eq. (2), we numerically solve it by the biconjugate gradient method Press-book . This numerical method has been successfully applied to obtain dark solitons or vertices in different models of Bose-Einstein condensates Winiecki-1999 ; Edmonds-2016 . By the method, when $\beta_{4}=1$, we obtain the numerical solutions of dark solitons under different $\beta_{2}$ and show them in Fig. 2. When $\beta_{2}=-1$, we find two kinds of fundamental SDSs, which has or does not have the total phase step (i.e., the key characteristic of dark solitons), and we show them at the top or bottom of Fig. 2 (a), respectively. Both of them have some stripes and oscillating tails, but their structures are different.
Later, we set their wave functions as the initial states to numerically integrate the model (2), by the split-step Fourier method Yang-book . The evolution results are shown as the insets in Fig. 2 (a). The SDS with the total phase step stably propagates for seven unit distance. At its final evolution stage, some new waves can be seen. Because the plane-wave background is unstable under the perturbation of dark solitons, these waves are generally induced by the modulation instability (MI), which will be analyzed in detail below. The similar result is also obtained for the SDS without the total phase step, which propagates stably for six unit distance.
@@
For some larger values of $\beta_{2}$, the results are shown in Figs. 2 (b), (c), (d), and (e). With the increase of $\beta_{2}$, the SDS has less stripes but can propagate for a longer distance, especially when $\beta_{2}>0$. Note that the SDS without the total phase step can be regarded as two separated dark solitons when $\beta_{2}=1.5$, as shown in Fig. 2 (e). It indicates that, when the SDS does not exist, the dark soliton without the total phase step does not exist.
III.2 Analysis of existence condition and oscillation frequency
@@
For the bright solitons in the model (2), the asymptotic analysis has been used to obtain their existence condition and oscillation characteristic Akhmediev-1994 ; Buryak-1995 ; Tam-2020 ; Zakharov-1998 . However, considering that the dark solitons have plane-wave backgrounds, the asymptotic analysis cannot be applied directly to them. As shown in Fig. 2, SDSs have the background waves, and meanwhile they have both of periodicity and localization. For such a nonlinear wave, the MLSA method can quantitatively give its dynamical characteristics and guide its generation Gao-2020 ; Gao-2021 . Thus, we will apply this method to analyze the existence condition and characteristics of fundamental SDSs. It starts with the ansatz solution of a perturbed plane wave,
@@
where the plane wave $\psi_{0}$ has the expression (3), and $u$ denotes a perturbation exerted on it. Substituting Eq. (4) into the model (2) and linearizing it about $u$, we can obtain
$i\frac{\partial u}{\partial z}-\frac{\beta_{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}+\frac{\beta_{4}}{24}\frac{\partial^{4}u}{\partial t^{4}}+\beta(u+u^{*})=0,$ (5)
where $u^{*}$ denotes the complex conjugation of $u$, and we have $\beta=a_{0}^{2}$. If the ansatz $u=e^{\lambda t}$ is applied to Eq. (5) according to Refs. Akhmediev-1994 ; Tam-2020 ; Zakharov-1998 , the term $u^{*}$ in Eq. (5) will be neglected and the results of dark solitons will be the same as the ones of bright solitons. It will cause that the derived existence region and oscillation frequency of SDSs don’t agree with the numerical results of Figs. 2 and 3 (b).
Therefore, to take the term $u^{*}$ into consideration, we assume the perturbation as the linear superposition of two conjugating components,
@@
where $A$ and $B$ are the amplitudes of the two components, and $p(t,z)$ is a complex function. This ansatz is a key setting of the MLSA method. Substituting Eq. (6) into the perturbation’s model (5) and only considering the parts of $e^{p(t,z)}$, one can obtain two linear equations of $A$ and $B^{*}$. They can be written as the following form,
$i\begin{bmatrix}-M+\beta&\beta\\ -\beta&M-\beta\end{bmatrix}\begin{bmatrix}A\\ B^{*}\end{bmatrix}=p_{z}\begin{bmatrix}A\\ B^{*}\end{bmatrix},$ (13)
@@
In the above expressions, $p_{t}$ and $p_{z}$ denote the partial derivative of $p(t,z)$ about $t$ and $z$, respectively. The function $p_{z}$ can be obtained by solving the eigenvalue of the left-hand matrix in Eq. (13):
$p_{z}=$$\pm\sqrt{M(2\beta-M)}.$ (15)
There are two modes, where the eigenvalues $p_{z}$ are the opposite of each other. The mode with a positive eigenvalue is focused on here.
@@
By this approximation, the function $p_{t}(t,z)$ can be successfully transformed into a complex constant, and therefore offers great convenience for our analysis. Moreover, this approximation can be conveniently applied to the function $p_{z}$. Considering that the cases of $t\rightarrow\pm\infty$ have the same results, we take the case of $t\rightarrow-\infty$ as an example and can obtain
$p_{z}^{(-)}=$ $\sqrt{M^{(-)}(2\beta-M^{(-)})},$ (18)
@@
It describes the propagation characteristics of the perturbation. To be specific, the perturbation’s propagation constant and growing rate relative to the plane wave are respectively
@@
From the insets of Fig. 2, we know that a SDS has a zero-velocity localized envelope and some zero-velocity stripes, which indicates $V_{\rm en}=0$ and $V_{\rm st}=0$. In Fig. 3 (a), when $\beta_{2}=-0.5$, $\beta_{4}=1$, and $\beta=1$, we show the curves of $V_{\rm en}=0$ and $V_{\rm st}=0$ in the $\omega$-$\eta$ plane. The coordinate of their intersection point $(\omega_{s},\eta_{s})$ is just the oscillation frequency and steepness of SDS, and the expressions are
$\omega_{s}={\rm Re}[\sqrt{\frac{-3\beta_{2}+2\sqrt{3\beta\beta_{4}}}{\beta_{4}}}],\;\eta_{s}={\rm Re}[\sqrt{\frac{3\beta_{2}+2\sqrt{3\beta\beta_{4}}}{\beta_{4}}}].$ (22)
@@
$\omega_{s}^{2}+\eta_{s}^{2}=\frac{4\sqrt{3\beta}}{\sqrt{\beta_{4}}},\quad\omega_{s}^{2}-\eta_{s}^{2}=-\frac{6\beta_{2}}{\beta_{4}}.$ (24)
@@
The former is independent of $\beta_{2}$, and the latter is independent of $\beta$. Thus, the point $(\omega_{s},\eta_{s})$ always locates on a circle when $\beta_{4}$ and $\beta$ are fixed, and it moves towards the positive axis of $\eta$ with $\beta_{2}$ increasing, as shown in Fig. 3 (a). More interestingly, this point is just an exceptional point of the eigenvalue $p_{z}$, which is related to the non-Hermiticity of the matrix in Eq. (13). By calculating the contour integral of its argument, one can obtain that its winding number is $\pi$ Ding-2022 . In eigenvalue problems, an exceptional point is a singularity where the two eigenvalues coalesce and each of them undergoes the transition from a real value to a complex. With the development of non-Hermitian physics, the exceptional point plays an important role in various systems, like optics, photonics, and Bose–Einstein condensates Miri-2019 ; El-Ganainy-2018 . However, for the relationship between SDS and exceptional point, a deeper understanding needs further investigations. Next, to verify the effectiveness of our prediction, the dependence of $\omega_{s}$ on $\beta_{2}$ is studied. When $\beta_{4}=1$ and $\beta=1$, we numerically measure the oscillation frequency of SDSs under different values of $\beta_{2}$ and show them by the red dots in Fig. 3 (b). One can see that $\omega_{s}$ gradually decreases into 0 with $\beta_{2}$ increasing. To obtain them, we measure the time interval between the first and second peaks on the right side of SDS as its time period $T_{s}$, and then the oscillation frequency can be calculated by $\omega_{s}=2\pi/T_{s}$, as shown in the upper row of Fig. 2 (a). On the other hand, we also show the analytic prediction from the MLSA method by the black curve. There are good agreements between the two results.
Figure 3: (Color online) (a) Phase of the propagating function $p_{z}(t,z)$ under different $\omega$ and $\eta$, when $\beta_{2}=-0.5$, $\beta_{4}=1$, and $\beta=1$. The black solid and dashed curves denote $V_{\rm en}=0$ and $V_{\rm st}=0$, respectively. The white point with the coordinate $(\omega_{s},\eta_{s})$ is an exceptional point of $\arg(p_{z})$ function. The yellow curve with an arrow is the trajectory of this point with $\beta_{2}$ increasing. (b) Dependence of SDS’s oscillation frequency $\omega_{s}$ on $\beta_{2}$, when $\beta_{4}=1$. The red dots and the black curve are respectively the results from numerical evolution and the MLSA method. (c) Dependence of the gain value $G$ on $\omega$, when $\beta_{2}=-1$ (upper) or $\beta_{2}=-0.3$ (lower). The strong or weak MI is defined by whether $\omega_{s}$ locates in the MI band or not. (d) Distribution of MI (modulation instability) and MS (modulation stability) on the $\beta_{2}$-$\beta_{4}$ plane. The color scale describes the value of $g(\omega_{s})$, i.e., Eq. (III.3).
@@
Figure 4: (Color online) The multi-dark-soliton bound states. The upper row is the distribution of their wave functions $\psi$ (red dashed curves) and amplitude $|\psi|$ (blue solid curves), when $\beta_{2}=0$, $\beta_{4}=1$, and $\beta=1$. The lower row is their corresponding amplitude evolution.
III.3 Analysis of propagation stability